Graph y=t+3sin⁡t and y=t on 0≤t≤2π. Where do the two graphs intersect on 0≤t≤2π?

Enter the exact answers in increasing order.

t = ?

t = ?

t = ?
Answer #1

The two graphs will intersect at those values of t for which t + 3sin(t) = t

.

Now t + 3sin(t) = t implies,

3sin(t) = 0

i.e. sin(t) = 0/3

i.e. sin(t) = 0

i.e. t = sin-1​​​​​​(0)

.

Now as 0 ≤ t ≤ 2π, so we have,

t = 0

t = π

t =